# There is No Quantum World

### v1

@@ Abstract
According to Aage Petersen [29, p. 12] who was Bohr’s assistant from 1952 to 1962, when Bohr was asked whether quantum mechanics could be understood as describing or mirroring an underlying quantum world, his response was
@@
This paper is divided into two parts. In the first part I present a neo-Bohrian interpretation of quantum mechanics—a view of quantum mechanics that accords with the core insights in Bohr’s thinking, with a twist that justifies the prefix ‘neo.’ The discussion here follows closely my earlier writing on this topic [13, 12]. The second part deals with the measurement problem, specifically how to deflate the measurement problem.
@@ 1 A Neo-Bohrian Interpretation of Quantum Mechanics
In 1935 Einstein, Podolsky, and Rosen [19] drew attention to the existence of non-separable or ‘entangled’ states of separated systems and exploited their correlations to argue for the incompleteness of quantum mechanics. John Bell re-examined the argument in 1964 in a seminal article [1] in which he proved that any deterministic extension of quantum mechanics with additional ‘hidden’ variables would necessarily introduce what Einstein [20] called ‘spooky action at a distance’ in the theory, in the sense that the outcome of a measurement could be influenced instantaneously by the setting of a remote measuring instrument. For Schrödinger, entanglement was ‘the characteristic trait of quantum mechanics, the one that enforces its entire departure from classical lines of thought’ [34, p. 555]. Following Bell’s proof, the probabilistic correlations of entangled quantum states are recognized as crucial to applications like quantum computation, quantum cryptography, and quantum communication.
@@
Classical theories are Boolean theories. To say that the algebra of observables of a quantum system is noncommutative is equivalent to saying that the sub-algebra of properties or propositions is non-Boolean. The non-Boolean algebra of quantum mechanics is formally a collection of Boolean algebras that are ‘intertwined,’ as Gleason put it [22, p. 886], in such a way that the whole collection can’t be embedded into a single inclusive Boolean algebra, so you can’t assign truth-values consistently to the propositions about observable values in all these Boolean algebras. In fact, there are finite sets of Boolean algebras in the collection of Boolean algebras of a quantum system that don’t fit together into a single Boolean algebra, unlike the corresponding collection for a classical system. Kochen and Specker, who first proved non-embeddability [25], call this structure a ‘partial Boolean algebra.’.
@@ 2 The Measurement Problem
There are different formulations of the measurement problem, but the classic formulation is von Neumann’s statement in [43, p. 351]. Von Neumann points to two fundamentally different ways in which the state of a quantum system can change: the change that occurs when a system undergoes measurement (he calls this Process 1), and the change when a system evolves dynamically under the action of an energy operator $H$ without measurement (Process 2). When a system evolves dynamically without being measured, the state $|\psi\rangle$ undergoes a continuous, reversible, unitary transformation in time $t$ to the state
@@
Schrödinger calls the measurement problem ‘the most difficult and most interesting point of the theory’ [32, p. 826]. Here’s how he put it
@@
For a physical system a measurement is a dynamical interaction with a second system that we characterize as a measuring instrument. The fact that we extract information about the measured system from the interaction has something to do with us, not the measured system. There should be no change in the nature of the dynamical behavior of the measured system just because we choose to regard the interaction as a measurement, so a special dynamics for measurement is unacceptable.
What would count as a solution to this problem? Firstly, we need to hold onto the fact that a measurement is a dynamical interaction between a measured system and a measuring instrument. Appeals to the consciousness of an observer who registers the measurement outcome can’t be invoked as a solution. The source of the problem is the noncommutativity or non-Booleanity of quantum mechanics, so we need to look to physical theory for a solution. Secondly, a salient feature of a quantum measurement, say of the spin of an electron, or the polarization of a photon, or the energy of a neutron, is that it involves a macroscopic setup in which we manipulate objects in our Boolean world to extract information about a microsystem, objects such as lasers, ion traps, optical tweezers, polarizing beamsplitters, photodetectors, interferometers, spectrometers, amplifiers, coincidence counters, and the like. The obvious solution would be a demonstration, if possible, that von Neumann’s Process 1 and Process 2 turn out to be equivalent, taking into account that any quantum measurement necessarily involves such systems.
Von Neumann’s work on infinite direct products of Hilbert spaces [41] (and later work with Murray on operator algebras [27]) is relevant here. I sketch the core idea following the account in Van Den Bossche and Grangier [35, 36, 37, 38].
@@
These considerations show that the measurement problem is resolved at the infinite limit. But infinity never comes. For macrosystems with very large but finite $N$, we can say only that interference between different macrostates is very small, less than any small number $\epsilon(N)$, and gets smaller and smaller, and harder and harder to detect, as the composite system includes more and more elementary systems. The approximate sectorization of Hilbert space as the number of elementary particles in a system increases without limit, and the consequent negligible interference between different macrostates, explains the classicality of the macroworld for all practical purposes—FAPP to use Bell’s acronym—and we can see why we experience the macroworld as classical or Boolean without the usual appeal to decoherence via loss of information to the environment. This is a nice result, but it does not show the equivalence between von Neumann’s Process 1 and Process 2.
Van Den Bossche and Grangier are clear about this in [35, p. 7]: ‘In principle the radical change of behaviour occurs in the $N\rightarrow\infty$ limit only ….’ But in [36, p. 5], they muddy the issue by saying that ‘even before reaching the $N\rightarrow\infty$ limit, the sectorisation behaviour sets in and converts the pure state in an effective mixed state,’ and ‘at some point, the number of electrons fed into the avalanche is so large that it results in a macroscopic change that cannot be ignored.’ (My emphasis.) The crucial words here are ‘effective’ and ‘at some point’ and ‘so large.’
@@
It is, however, reducible in a different sense. A measurement process can be analyzed to any level of precision in a quantum theoretical treatment, but some ‘ultimate’ macroscopic detecting instrument must always be left out of the quantum mechanical analysis as a classical/Boolean reference system for the Born probabilities of the quantum state, which are interpreted as referring to possible measurement outcomes indicated by the ‘pointer readings’ of the detector, one of which actually occurs. The detector itself could, in principle, be analyzed as a finite quantum mechanical system, with the Born probabilities then referring to a super-detector. The movability of the “Heisenberg cut’ between the system described in terms of quantum mechanics and the macrosystem acting as the detecting instrument is what justifies the account as complete and objective, in the sense that anything relevant to the measurement can potentially be included in the analysis. As Peres put it [28], ‘although it can describe anything, a quantum description cannot include everything.’ On this neo-Bohrian interpretation of quantum mechanics the measurement problem is not resolved but deflated as a feature of non-Booleanity.
@@ Acknowledgements
Thanks to Michel Janssen, Allen Stairs, and especially Michael Cuffaro, for insightful and extremely helpful comments.

### v2

# Jeffrey Bub

@@ Abstract
According to Aage Petersen [32, p. 12] who was Bohr’s assistant from 1952 to 1962, when Bohr was asked whether quantum mechanics could be understood as describing or mirroring an underlying quantum world, his response was
@@
This paper is divided into two parts. In the first part I present a neo-Bohrian interpretation of quantum mechanics—a view of quantum mechanics that accords with the core insights in Bohr’s thinking, with a twist that justifies the prefix ‘neo.’ The discussion here follows closely my earlier writing on this topic [13, 12]. The second part deals with the measurement problem, specifically how to put away the measurement problem.
@@ 1 A Neo-Bohrian Interpretation of Quantum Mechanics
In 1935 Einstein, Podolsky, and Rosen [20] drew attention to the existence of non-separable or ‘entangled’ states of separated systems and exploited their correlations to argue for the incompleteness of quantum mechanics. John Bell re-examined the argument in 1964 in a seminal article [1] in which he proved that any deterministic extension of quantum mechanics with additional ‘hidden’ variables would necessarily introduce what Einstein [21] called ‘spooky action at a distance’ in the theory, in the sense that the outcome of a measurement could be influenced instantaneously by the setting of a remote measuring instrument. For Schrödinger, entanglement was ‘the characteristic trait of quantum mechanics, the one that enforces its entire departure from classical lines of thought’ [37, p. 555]. Following Bell’s proof, the probabilistic correlations of entangled quantum states are recognized as crucial to applications like quantum computation, quantum cryptography, and quantum communication.
@@
Classical theories are Boolean theories. To say that the algebra of observables of a quantum system is noncommutative is equivalent to saying that the sub-algebra of properties or propositions is non-Boolean. The non-Boolean algebra of quantum mechanics is formally a collection of Boolean algebras that are ‘intertwined,’ as Gleason put it [23, p. 886], in such a way that the whole collection can’t be embedded into a single inclusive Boolean algebra, so you can’t assign truth-values consistently to the propositions about observable values in all these Boolean algebras. In fact, there are finite sets of Boolean algebras in the collection of Boolean algebras of a quantum system that don’t fit together into a single Boolean algebra, unlike the corresponding collection for a classical system. Kochen and Specker, who first proved non-embeddability [26], call this structure a ‘partial Boolean algebra.’.
@@ 2 The Measurement Problem
There are different formulations of the measurement problem, but the classic formulation is von Neumann’s statement in [46, p. 351]. Von Neumann points to two fundamentally different ways in which the state of a quantum system can change: the change that occurs when a system undergoes measurement (he calls this Process 1), and the change when a system evolves dynamically under the action of an energy operator $H$ without measurement (Process 2). When a system evolves dynamically without being measured, the state $|\psi\rangle$ undergoes a continuous, reversible, unitary transformation in time $t$ to the state
@@
For a physical system a measurement is a dynamical interaction with a second system that we characterize as a measuring instrument. The fact that we extract information about the measured system from the interaction has something to do with us, not the measured system. There should be no change in the nature of the dynamical behavior of the measured system just because we choose to regard the interaction as a measurement. The problem is that quantum mechanics requires a dual dynamics, and this is unacceptable without further explanation.
Schrödinger calls the measurement problem ‘the most difficult and most interesting point of the theory’ [35, p. 826]. Here’s how he put it
@@
What would count as a solution to this problem? Firstly, we need to hold onto the fact that a measurement is a dynamical interaction between a measured system and a measuring instrument. Appeals to the consciousness of an observer who registers the measurement outcome can’t be invoked as a solution. The source of the problem is the noncommutativity or non-Booleanity of quantum mechanics, so we need to look to physical theory for a solution. Secondly, a salient feature of a quantum measurement, say of the spin of an electron, or the polarization of a photon, or the energy of a neutron, is that it involves a macroscopic setup in which we manipulate objects in our Boolean world to extract information about a microsystem, objects such as lasers, ion traps, optical tweezers, polarizing beamsplitters, photodetectors, interferometers, spectrometers, amplifiers, coincidence counters, and the like. The obvious solution would be a demonstration, if possible, that von Neumann’s Process 1 can be reduced to the unitary dynamics of Process 2, taking into account that any quantum measurement necessarily involves such systems.
Von Neumann’s work on infinite direct products of Hilbert spaces [44] (and later work with Murray on operator algebras [30]) is relevant here. I sketch the core idea following the account in Van Den Bossche and Grangier [38, 39, 40, 41]. See also Landsman [27, 28] and Butterfield [15].
@@
These considerations show that the measurement problem is resolved at the infinite limit. But infinity never comes. For macrosystems with very large but finite $N$, we can say only that interference between different macrostates is very small, less than any small number $\epsilon(N)$, and gets smaller and smaller, and harder and harder to detect, as the composite system includes more and more elementary systems. The approximate sectorization of Hilbert space as the number of elementary particles in a system increases without limit, and the consequent negligible interference between different macrostates, explains the classicality of the macroworld for all practical purposes—FAPP to use Bell’s acronym—and we can see why we experience the macroworld as classical or Boolean without the usual appeal to decoherence via loss of information to the environment. This is a nice result, but it does not explain the occurrence of a definite outcome in a quantum measurement process, so it does not explain how von Neumann’s Process 1 is reducible to Process 2.
Van Den Bossche and Grangier are clear about this in [38, p. 7]: ‘In principle the radical change of behaviour occurs in the $N\rightarrow\infty$ limit only ….’ But in [39, p. 5], they muddy the issue by saying that ‘even before reaching the $N\rightarrow\infty$ limit, the sectorisation behaviour sets in and converts the pure state in an effective mixed state,’ and ‘at some point, the number of electrons fed into the avalanche is so large that it results in a macroscopic change that cannot be ignored.’ (My emphasis.) The crucial words here are ‘effective’ and ‘at some point’ and ‘so large.’
@@
Where does this leave things? We know that something happens in a quantum measurement—we see an event occurring, the registration of a definite outcome at the macrolevel. A non-Boolean theory cannot explain how this happens. Yes, interference will be smaller than any designated number $\epsilon(N)$ for large $N$, but for any $N$, no matter how large, there is some interference experiment that could in principle be performed (granted, with practically impossible difficulty) that would reveal interference inconsistent with the measured observable actually taking a definite value—one could derive a contradiction from the assumption that the measured observable takes a definite value and the existence of interference, however negligible.
So, the occurrence of a definite measurement outcome event must be outside the quantum description. A measurement process can be analyzed to any level of precision in a quantum theoretical treatment, but some ‘ultimate’ macroscopic detecting instrument must always be left out of the non-Boolean analysis as a classical/Boolean reference system for the Born probabilities of the quantum state, which are interpreted as referring to possible measurement outcomes indicated by the ‘pointer readings’ of the detector, one of which actually occurs. The detector itself could, in principle, be analyzed as a finite quantum mechanical system, with the Born probabilities then referring to a classical/Boolean super-detector. While the requirement to leave some part of an experimental set-up outside the quantum description cannot be avoided, the movability of the “Heisenberg cut’ between the system described quantum mechanically and the macrosystem acting as the detecting instrument is what justifies the account as complete and objective, in the sense that anything relevant to the measurement can potentially be included in the analysis. As Peres put it [31], ‘although it can describe anything, a quantum description cannot include everything.’ In this sense of ‘successive’ reduction, in which a larger and larger part of an experimental set-up can be treated quantum mechanically, without any limit in principle, the Boolean macroworld is reducible to the non-Boolean quantum level. On this neo-Bohrian interpretation of quantum mechanics the measurement problem is not resolved but deflated as a feature of non-Booleanity.
@@ Acknowledgements
Thanks to Michel Janssen, Allen Stairs, and especially Michael Cuffaro and Jeremy Butterfield, for insightful and extremely helpful comments.